how many gallons of an 80% solution must be mixed with 90 gallons of 15% solution to make a 70% solution?
Answer by stanbon(75887)

how many gallons of an 80% solution must be mixed with 90 gallons of 15% solution to make a 70% solution?
---------------------
Equation:
active + actve = active
0.15*90 + 0.80x = 0.70(90+x)
---------
15*90 + 80x = 70*90 +70x
10x = 55*90
x = 9*55 = 495
--------------
Cheers,
Stan H.
==============